# Frontiers of Generative AI for Network Optimization: Theories, Limits, and Visions

### v1

@@ VI Future Directions
Despite the impressive performance of GenAI in network optimization and related AI domains, debates over its internal mechanisms remain unresolved. In recent years, both significant benchmark improvements driven by new GenAI techniques and critical analyses exposing their limitations have emerged in parallel. Even within the AI community, opinions diverge: some believe that the continuous breakthroughs of LPTMs and GDMs across various metrics indicate their potential to approach human-level intelligence; others argue that these gains are merely incremental advances fueled by data scale and engineering tricks, rather than evidence of genuine progress toward human-like reasoning. In reality, both perspectives largely rely on empirical results obtained through specific testing, while deep theoretical understanding remains scarce, mainly because many phenomena observed in LPTMs and GDMs are still poorly understood.
In light of this, we argue that both sides of the evidence should be considered. On one hand, we should leverage current, effective techniques to continue improving performance metrics; on the other, we must not let these successes overshadow the models’ limitations. It is essential to continue identifying flaws and exploring new directions. This section explores potential future directions for GenAI in network optimization, from both theoretical and implementation perspectives.
@@ VI-A1 Disentangling the Link Between Generation and Optimization
For example, it is not yet well understood whether common optimization problem classes (e.g., convex optimization, mixed-integer programming, combinatorial optimization) exhibit inherent compatibility with typical generative distributions (e.g., Gaussian, Poisson, or multi-modal distributions). Clarifying these relationships would help define the expected target distribution when applying GenAI to network optimization problems, thereby enabling better estimates of learning difficulty and theoretical performance limits. This is especially crucial when dealing with long-tailed distributions, which are notoriously challenging for GenAI models. A deeper theoretical understanding in this direction can improve both the design and deployment of GenAI-based solvers under various problem regimes.
@@ VI-A2 Decomposing Network Optimization Problems
A classical example lies in transforming non-convex problems into convex ones through relaxation techniques. In this vein, GenAI models could be developed as specialized solvers targeting different types of simpler sub-problems, with their outputs fused to form a high-quality overall solution. This modular approach reduces both the complexity of designing a single end-to-end GenAI optimizer and improves generalizability across tasks.
@@ VI-B1 Rule-Guided Generation
Recent studies have consistently shown that current GenAI models—particularly LPTMs—do not exhibit human-like capabilities in conceptual understanding or rule capturing [165, 166]. Although reinforcement learning has been used to enhance reasoning, these improvements mainly reflect behavioral preference tuning rather than genuine comprehension of underlying logic [47]. Despite growing skepticism, no fundamentally new paradigm has yet emerged to replace autoregressive generation or sampling-based methods.
@@ VI-B2 Collaboration Beyond Isolated Decision-Making
Recent advances in GenAI have demonstrated significant improvements in mathematical reasoning, formal modeling, and symbolic proof generation [174, 175, 76, 75, 74, 49]. While the underlying mechanisms behind these capabilities remain incompletely understood, engineering outcomes are increasingly competitive in specialized domains. Given GenAI’s inherent uncertainty, imprecision, and reliance on training data distributions, many of these works have adopted modular architectures, decomposing the complex task of mathematical problem solving into multiple cooperative components.
In the field of network optimization, which demands structure and precision, early studies such as [77, 82] offer valuable insights. For instance, [77] decomposes an optimization problem into five components, including original context, physical parameters, optimization variables, objective functions, and constraints, using natural language prompts for formalization. Meanwhile, [82] leverages LLMs to orchestrate the interaction with external solvers. These designs demonstrate the value of GenAI as a component within a broader, orchestrated system.
Building on this perspective, we argue that true reliability and performance in network optimization may be achieved not through standalone GenAI models but through synergistic collaboration between GenAI, traditional numerical solvers, conventional neural networks, and human expertise. We envisage a five-module cooperative framework: a Modeler that interprets user intent and formulates formal optimization problems; an Optimizer that performs core numerical or data-driven solving; an Invoker that integrates external toolchains or simulation engines; a Verifier that evaluates solution validity and feasibility; and a Human Expert who intervenes when model limitations are reached, providing feedback and corrective strategy. Such a collaborative system mitigates the intrinsic limitations of GenAI by distributing cognitive load and leveraging complementary strengths across components. Moreover, it aligns well with distributed or decentralized deployment environments common in real-world network systems [176]. Importantly, this architecture fosters the potential for autonomous evolution, allowing GenAI agents to learn from human interventions over time, progressively enhancing their capabilities in dynamic, complex environments.
@@ VI-B3 Virtualization-Enabled Evolution
Beyond the development and deployment of GenAI models, it is equally important to consider how these models evolve alongside dynamic network environments. Traditional static deployment strategies fall short in long-term performance maintenance, especially in systems that require adaptability to fluctuating traffic, topology changes, and real-time decision-making. We therefore advocate for a virtualized evolution infrastructure, enabling GenAI models to be continuously refined, evaluated, and validated in a controlled, cost-efficient environment.
We envision a virtual evolution platform that mirrors real-world network dynamics and allows GenAI agents to simulate their deployment scenarios before going live, like digital twins. This environment should be tightly synchronized with real-time data flows from the physical network, offering rich contextual information for pre-deployment testing. Recently, standardization institutes such as european telecommunications standards institute (ETSI) and next generation mobile networks alliance (NGMN) have begun integrating sandboxing, public memory, and other virtualization primitives into their Agent-based AI-Core architectures [177, 178], recognizing the importance of controlled environments for AI development and orchestration. This approach enables GenAI models to interact with realistic network conditions, gather feedback, and adjust behavior without incurring the cost or risk of live network disruption. It also allows developers to conduct large-scale experiments and monitor model behaviors across diverse conditions, fostering robust generalization and reliability.
While existing digital twin frameworks have been applied in networking contexts, most current efforts focus on optimizing algorithmic performance within the virtual layer, rather than designing the virtualization infrastructure needed for GenAI-driven evolution. We believe that achieving a next-generation, evolution-capable AI network requires significant advancements in three areas, including efficient collection and representation of high-dimensional, heterogeneous environment data; memory and retrieval mechanisms for distilled knowledge, enabling cross-task transfer and generalization; and fine-grained temporal and spatial simulation of network dynamics to improve model adaptability and resilience.
VII Existing Benchmarks and Datasets
TABLE VIII: Relevant benchmarks and datasets, with the relevance to network optimization.
{tblr}
colspec=X[0.15,l] X[0.15,c] X[0.08,c] X[0.6,c], cells = c, hlines, vlines, rowsep = 2pt, row1 = 0ex, cell21 = r=3, cell51 = r=2, cell71 = r=3, cell101 = r=3, Type & Reference Relevance Content
@@ VII Existing Benchmarks and Datasets
To complement the theoretical and methodological discussions above, we compile and categorize a diverse set of recent benchmarks and datasets relevant to network optimization in Table VIII. These resources are organized into four categories: (1) mathematical question-answering (QA) benchmarks that assess symbolic reasoning and mathematical problem-solving ability; (2) telecom-specific QA datasets tailored for LLM-based understanding of domain-specific language and protocols; (3) network optimization benchmarks or task datasets, which directly represent optimization problem instances or user intent in natural language; and (4) related upstream and downstream task datasets, which contribute to modeling, perception, and interaction in network environments, though they do not directly pose optimization problems.
@@
Collectively, these datasets underscore the growing ecosystem supporting GenAI-driven network optimization. They enable more rigorous benchmarking, facilitate model transferability, and reveal application gaps. However, challenges remain. General-purpose benchmarks often overlook optimization-specific criteria such as feasibility, constraint satisfaction, and convergence behavior. Looking forward, there is a clear demand for more comprehensive and standardized benchmarks that can evaluate GenAI models in both problem understanding and high-precision decision-making. Future datasets should integrate real-world conditions, include structured constraint annotations, and support multi-stage inference pipelines.
@@ VIII Conclusion
Generative models hold transformative potential for the field of network optimization, offering more efficient and higher-performing approaches to predicting and optimizing communication systems. This survey has focused on the two most prominent categories of GenAI models—GDMs and LPTMs—to systematically review the latest theoretical and implementation progress. We began by examining relevant developments in the AI community that have inspired GenAI applications in networking, then categorized existing GenAI efforts within the network domain. Drawing on the two predominant modeling paradigms in network optimization (one-shot optimization and MDP) formulations, we classified and analyzed the roles and capabilities of GDMs and LPTMs across these contexts.
We have revealed that the foundational theories and initial implementations of GenAI in network optimization are already in place. Both AI and networking communities are actively pursuing all-in-one, pre-trained models aimed at general-purpose applications. However, we also observed several unresolved challenges. GenAI’s inherently probabilistic nature continues to hinder precise problem-solving, particularly when strict constraints must be satisfied. LPTMs face additional limitations, such as excessive model size and computational overhead.
Beyond reviewing existing work, we further have integrated and discuss recent critical perspectives from both the AI and networking communities. These contributions highlight fundamental limitations that are often overshadowed by the impressive surface-level performance of GenAI. Notably, we have presented generalization error bounds between generated and optimal solutions (in one-shot settings) and between generated and optimal policies (in MDP settings), providing theoretical guidance for GenAI model development. We then discussed GenAI’s strengths and weaknesses in network optimization and propose future research directions from both theoretical and implementation perspectives, emphasizing the importance of bridging the generation-optimization gap and incorporating rule-guided generation ideas.

### v2

@@ VI Future Directions
Despite the impressive performance of GenAI in network optimization and related AI domains, debates over its internal mechanisms remain unresolved. In recent years, both significant benchmark improvements driven by new GenAI techniques and critical analyses exposing their limitations have emerged in parallel. Even within the AI community, opinions diverge: some believe that the continuous breakthroughs of LPTMs and GDMs across various metrics indicate their potential to approach human-level intelligence; others argue that these gains are merely incremental advances fueled by data scale and engineering tricks, rather than evidence of genuine progress toward human-like reasoning. In reality, both perspectives largely rely on empirical results from specific testing, while deep theoretical understanding remains scarce, mainly because many phenomena observed in LPTMs and GDMs remain poorly understood. In the summary of cutting-edge works presented in Table IX, we illustrate the diverse attitudes of current researchers in this area.
In light of this, we argue that both sides of the evidence should be considered. On the one hand, we should leverage current, effective techniques to continue improving performance metrics. On the other hand, we must not let these successes overshadow the models’ limitations. It is essential to continue identifying flaws and exploring new directions. This section explores potential future directions for GenAI in network optimization, from both theoretical and implementation perspectives. These directions not only reflect the potential application of GenAI in network optimization, but also point out some specific problems that may be helped to be solved in the current network optimization domain, such as explainability of the principles, diversity of solvable problems, and reliability of the solutions.
@@ VI-A1 Disentangling the Link Between Generation and Optimization
For example, it is not yet well understood whether common optimization problem classes (e.g., convex optimization, mixed-integer programming, combinatorial optimization) are inherently compatible with typical generative distributions (e.g., Gaussian, Poisson, or multimodal distributions). Clarifying these relationships would help define the expected target distribution when applying GenAI to network optimization problems, thereby enabling better estimates of learning difficulty and theoretical performance limits. This is especially crucial when dealing with long-tailed distributions, which are notoriously challenging for GenAI models. A deeper theoretical understanding in this direction can improve both the design and deployment of GenAI-based solvers under various problem regimes.
@@ VI-A2 Decomposing Network Optimization Problems
A classical example is transforming non-convex problems into convex ones using relaxation techniques. In this vein, GenAI models could be developed as specialized solvers targeting different types of simpler sub-problems, with their outputs fused to form a high-quality overall solution. This modular approach reduces both the complexity of designing a single end-to-end GenAI optimizer and improves generalizability across tasks.
@@
Figure 6: Outlook on the applications of GenAI in the AI-Core and next-generation networking.
@@ VI-B1 Rule-Guided Generation
Recent studies have consistently shown that current GenAI models—particularly LPTMs—do not exhibit human-like capabilities in conceptual understanding or rule capturing [163]. Although reinforcement learning has been used to enhance reasoning, these improvements mainly reflect behavioral preference tuning rather than genuine comprehension of underlying logic. [47] investigates whether RL with verifiable rewards (RLVR) truly enhances LLM reasoning beyond base models, finding that current RLVR methods do not elicit fundamentally new reasoning patterns but are bounded by the base model’s capabilities. Despite growing skepticism, no fundamentally new paradigm has yet emerged to replace autoregressive generation or sampling-based methods.
@@
Figure 7: Current advantages, limitations, and the corresponding relevant works of GenAI for network optimization.
@@ VI-B2 Collaboration Beyond Isolated Decision-Making
Recent advances in GenAI have demonstrated significant improvements in mathematical reasoning, formal modeling, and symbolic proof generation [173, 174, 76, 75, 74, 49]. While the underlying mechanisms behind these capabilities remain incompletely understood, engineering outcomes are increasingly competitive in specialized domains. Given GenAI’s inherent uncertainty, imprecision, and reliance on training data distributions, many of these works have adopted modular architectures that decompose the complex task of mathematical problem-solving into multiple cooperative components.
In the field of network optimization, which demands structure and precision, early studies such as [77, 54] offer valuable insights. For instance, [77] decomposes an optimization problem into five components, including original context, physical parameters, optimization variables, objective functions, and constraints, using natural language prompts to formalize it. Meanwhile, [54] leverages LLMs to orchestrate the interaction with external solvers. These designs demonstrate the value of GenAI as a component within a broader, orchestrated system.
Building on this perspective, we argue that true reliability and performance in network optimization may be achieved not through standalone GenAI models but through synergistic collaboration between GenAI, traditional numerical solvers, conventional neural networks, and human expertise. We envisage a five-module cooperative framework: a Modeler that interprets user intent and formulates formal optimization problems; an Optimizer that performs core numerical or data-driven solving; an Invoker that integrates external toolchains or simulation engines; a Verifier that evaluates solution validity and feasibility; and a Human Expert who intervenes when model limitations are reached, providing feedback and a corrective strategy. Such a collaborative system mitigates the intrinsic limitations of GenAI by distributing cognitive load and leveraging complementary strengths across components. Moreover, it aligns well with distributed or decentralized deployment environments common in real-world network systems [175]. Importantly, this architecture enables autonomous evolution, allowing GenAI agents to learn from human interventions over time and progressively enhance their capabilities in dynamic, complex environments.
@@ VI-B3 Virtualization-Enabled Evolution
Beyond the development and deployment of GenAI models, it is equally important to consider how these models evolve alongside dynamic network environments. Traditional static deployment strategies fall short in long-term performance maintenance, especially in systems that require adaptability to fluctuating traffic, topology changes, and real-time decision-making. We therefore advocate for a virtualized evolution infrastructure that enables GenAI models to be continuously refined, evaluated, and validated in a controlled, cost-efficient environment.
We envision a virtual evolution platform that mirrors real-world network dynamics and enables GenAI agents to simulate their deployment scenarios before going live, akin to digital twins. This environment should be tightly synchronized with real-time data flows from the physical network, offering rich contextual information for pre-deployment testing. Recently, standardization institutes such as European Telecommunications Standards Institute (ETSI) and Next Generation Mobile Networks Alliance (NGMN) have begun integrating sandboxing, public memory, and other virtualization primitives into their Agent-based AI-Core architectures [176, 177], recognizing the importance of controlled environments for AI development and orchestration. This approach enables GenAI models to interact with realistic network conditions, gather feedback, and adjust behavior without incurring the cost or risk of live network disruption. It also allows developers to conduct large-scale experiments and monitor model behaviors across diverse conditions, fostering robust generalization and reliability.
TABLE X: Relevant benchmarks and datasets, with the relevance to network optimization.
Type Reference Relevance Content
@@
While existing digital twin frameworks have been applied in networking contexts, most current efforts focus on optimizing algorithmic performance within the virtual layer, rather than designing the virtualization infrastructure needed for GenAI-driven evolution. We believe that achieving a next-generation, evolution-capable AI network requires significant advancements in three areas, including efficient collection and representation of high-dimensional, heterogeneous environment data; memory and retrieval mechanisms for distilled knowledge, enabling cross-task transfer and generalization; and fine-grained temporal and spatial simulation of network dynamics to improve model adaptability and resilience. Fig. 6 illustrates our vision for the application scenarios and functional roles of GenAI within the ETSI AI-Core architecture [176].
VI-C Empirical Suggestions
In the context of the evolving landscape of GenAI for network optimization and related fields, we summarize the key milestones and latest advancements. As illustrated in Fig. 7, we present a synthesized view of the technology’s inherent performance advantages and critical limitations. Specifically, positive experiences, such as superiority to discriminative models, robustness to data quality, enhanced cross-task generalization, mathematical reasoning, and proficient task intent extraction, are often underpinned and validated by pioneering works [106, 109, 92, 45, 54]. Conversely, several critical limitations have emerged, including limited self-evolution/correction, weak practical cross-task performance compounded by insufficient theoretical basis, output uncertainty and reliability risks related to solution validity, and limited conceptual comprehension. These weaknesses have been identified by critical studies [44, 48, 103, 165] which often point toward viable future research directions. By distilling these positive and negative empirical experiences, we aim to provide readers and practitioners with direct guidance for both effective deployment and focused future investigation.
Based on our past research experience and the conclusions drawn from related work, we propose three valuable recommendations for effective implementation and the avoidance of common pitfalls. First, data diversity is key to mitigating long-tail bias. The training dataset must be sufficiently diverse. Otherwise, the model will suffer from long-tail distribution bias, causing the generator to disproportionately produce the most frequent solution category during inference. Second, structural alignment in model selection. Model architectures should align with the problem type. Specifically, structured or sequential optimization problems may find GCNs and Transformer architectures more suitable. Third, handling GDM loss convergence anomalies. In GDMs for network optimization, an occasional observation is a rapid decrease in loss followed by a sharp rise and a prolonged plateau. This behavior often stems from the task’s complexity, creating large high-loss planes in the parameter space, and the issue is typically resolved by simply restarting the training process.
VII Existing Benchmarks and Datasets
To complement the theoretical and methodological discussions above, we compile and categorize a diverse set of recent benchmarks and datasets relevant to network optimization in Table X. These resources are organized into four categories: (1) mathematical question-answering (QA) benchmarks that assess symbolic reasoning and mathematical problem-solving ability by accuracy; (2) telecom-specific QA datasets that assess LLM-based understanding of domain-specific language and protocols by selection accuracy and contextual accuracy; (3) network optimization benchmarks or task datasets that evaluate the model’s ability to interpret and solve optimization-related problems by accuracy, optimality gap and feasibility rate; and (4) upstream and downstream task datasets that measure modeling, perception, and interaction capabilities in network environments through multiple metrics, including accuracy, efficiency, and quality.
@@ VII Existing Benchmarks and Datasets
Collectively, these datasets underscore the growing ecosystem supporting GenAI-driven network optimization. They enable more rigorous benchmarking, facilitate model transferability, and reveal application gaps. However, challenges remain. General-purpose benchmarks often overlook optimization-specific criteria such as feasibility, constraint satisfaction, and convergence behavior. Looking ahead, there is clear demand for more comprehensive, standardized benchmarks to evaluate GenAI models across problem understanding and high-precision decision-making. Future datasets should integrate real-world conditions, include structured constraint annotations, and support multi-stage inference pipelines.
@@ VIII Conclusion
Generative models hold transformative potential for network optimization, offering more efficient, higher-performing approaches to predicting and optimizing communication systems. This survey has explored the theories, limitations and visions of GenAI in the field of network optimization. We began by examining relevant developments in the AI community that have inspired GenAI applications in networking, and then we categorized existing GenAI efforts within the network domain. Building on this foundation, we comprehensively reviewed the latest theoretical and implementation advancements of the two most representative GenAI models—GDMs and LPTMs—and further analyzed their roles and capabilities.
We have revealed that the foundational theories and initial implementations of GenAI in network optimization are already in place. Both the AI and networking communities are actively pursuing all-in-one, pre-trained models for general-purpose applications. However, we identified unresolved challenges, including limitations in LPTM’s mathematical reasoning and conceptual understanding, limitations of rule learning by GDMs, and difficulties meeting real-time requirements for network optimization.
Beyond reviewing existing work, we further have integrated and discuss recent critical perspectives from both the AI and networking communities. These contributions highlight fundamental limitations that are often overshadowed by the impressive surface-level performance of GenAI. Notably, we have presented generalization error bounds between generated and optimal solutions (in one-shot settings) and between generated and optimal policies (in MDP settings), providing theoretical guidance for GenAI model development. We then discussed GenAI’s strengths and weaknesses in network optimization and proposed future research directions from both theoretical and implementation perspectives, emphasizing the importance of bridging the generation-optimization gap and incorporating rule-guided generation ideas. We also provide valuable positive and negative research experiences, hoping to guide other researchers.